# Expanding the Resolution Boundary of Outcome-Based Imperfect-Recall Abstraction in Games with Ordered Signals

### v1

@@ 9 Experiment and Result
Experiments were conducted on Numeral211 Hold’em, a simplified version of two-player limit Texas Hold’em, which is divided into three stages; specific rules are detailed in Appendix D.2. In Numeral211 Hold’em, the quantity of abstracted signal infosets identified by each method at different stages is presented in Table 3. The performance of 0-Recall Outcome Isomorphism (i.e., Potential Outcome Isomorphism, POI), 2-Recall Outcome Isomorphism (2-ROI), and Lossless Isomorphism (LI) was evaluated in following experiments.
Stage 00-ROI $1$-ROI $2$-ROI LI
@@ 9.1 Asymmetric Abstraction in CFR Training
Our fisrt investigation focuses on the one-player-abstraction-perspective. Given a signal abstraction $\tilde{\alpha}^{\varTheta}=(\tilde{\alpha}^{\varTheta}_{1},\tilde{\alpha}^{\varTheta}_{2})$, we construct scenarios where player 1 employs the signal abstraction, while player 2 does not, resulting in the signal abstracted game $\tilde{\Gamma}^{1}=\tilde{\Gamma}^{(\tilde{\alpha}^{\varTheta}_{1},\varTheta_{2})}$. Conversely, we adjust the setup such that player 1 does not use abstraction, whereas player 2 does, thereby creating another signal abstracted game $\tilde{\Gamma}^{2}=\tilde{\Gamma}^{(\varTheta_{1},\tilde{\alpha}^{\varTheta}_{2})}$. Subsequently, by applying the chance-sampled counterfactual regret minimization (CSMCCFR) (Zinkevich et al., 2007; Lanctot et al., 2009), we iteratively solve for $\epsilon$-Nash equilibrium and map the resulting strategies back to the original game $\tilde{\Gamma}$ to obtain $\sigma^{1,*}=(\sigma^{1,*}_{1},\sigma^{1,*}_{2})$ and $\sigma^{2,*}=(\sigma^{2,*}_{1},\sigma^{2,*}_{2})$. Ultimately, by amalgamating these strategies into $\sigma^{*}=(\sigma^{1,*}_{1},\sigma^{2,*}_{2})$, which we called asymmetric abstraction Nash strategy, we adopt the $\sigma^{*}$ strategy in real-game scenarios as a means to counter strategies available in a non-abstracted setting from an abstracted viewpoint. The exploitability of $\sigma$ is then analyzed in terms of milli blinds per game (mb/g). The experimental results are depicted in Figure 6, where (a) showcases the results plotted on a linear scale of exploitability, while (b) presents the same results on a logarithmic scale.
@@ 9.2 Symmetric Abstraction in CFR Self-Play Training
Subsequently, we shift our focus to the setting of symmetric abstraction, which more closely aligns with the practical scenarios of AI training. Given a specific signal abstraction $\tilde{\alpha}^{\varTheta}$, we construct the corresponding signal abstracted game $\tilde{\Gamma}^{\tilde{\alpha}^{\varTheta}}$. Utilizing the CSMCCFR algorithm, we iteratively solve for the $\epsilon$-Nash equilibrium and map the resulting strategy back to the original game $\tilde{\Gamma}$, obtaining $\sigma^{*}$, and then calculate its exploitability. The experimental results are presented in Figure 7.
@@ Appendix C Perfect-Recall Games
    In any path of $\Gamma$, no two nodes within the same information set $I\in\mathcal{I}_{i}$ of player $i$ can occur simultaneously. In this case, for any node $x$ satisfying $I\sqsubseteq x$, $I[x]$ denotes the node in the path from $x^{o}$ to $x$ that belongs to $I$.
@@ D.1 Leduc Hold’em
Leduc Hold’em is a poker variant designed for two players, conceived primarily as a platform for computer game-playing research and introduced by Waugh et al. (2009a). It was engineered to possess a strategic depth akin to Texas Hold’em while maintaining a manageable scale suitable for the development of intelligent gameplay strategies.
@@
    First Betting Stage: Following the distribution of hole cards, a stage of betting occurs. Players can choose to check or bet, with the bet size set at two chips.
@@
    Flop: After the initial betting stage, a single community card, termed the flop, is revealed from the deck.
@@
    Second Betting Stage: Another stage of betting takes place after the flop, with the bet size increasing to four chips.
@@
    Betting Options: Throughout the game, players have options to fold, call, or raise. Each betting stage permits only one bet and one raise per player, with fixed bet sizes of two chips in the first stage and four chips in the second.
@@ D.2 Numerall211 Hold’em
    First Betting Stage: Following the distribution of hole cards, a stage of betting occurs. Players can choose to check or bet, with the bet size set at 10 chips.
@@
    Flop: After the initial betting stage, a single community card, termed the flop, is revealed from the deck.
@@
    Second Betting Stage: Another stage of betting takes place after the flop, with the bet size increasing to 20 chips.
@@
    Turn: After the Second betting stage, another community card, termed the turn, is revealed from the deck.
@@
    Third Betting Stage: Another stage of betting takes place after the turn, with the bet size still set at 20 chips.
@@
    Betting Options: Throughout the game, players have options to fold, call, or raise. In each betting stage, the total sum of bets and raises is limited to a maximum of 4, with fixed bet sizes of 10 chips in the first stage and 20 chips in the last two betting stages.
@@
1 Straight flush 0.00321
3 of cards with consecutive rank and same suit. Ties are broken by highest card.
@@
2 Three of a kind 0.01587
3 of cards with the same rank. Ties are broken by the card’s rank.
@@
3 Straight 0.04347
3 of cards with consecutive rank. Ties are broken by the highest card rank.
@@
4 Flush 0.15799
3 of cards with the same suit. Ties are broken by the highest card rank, then second highest card rank, then third highest card rank.
@@
5 Pair 0.34065
2 of cards with the same rank. Ties are broken by the rank of the pair, then by the rank of the third card.
@@
6 High card 0.43881
None of the above. Ties are broken by comparing the highest ranked card, then the second highest ranked card, and then the third highest ranked card
@@ Appendix E A Transition from Games with Ordered Signals to Imperfect-Information Games
The games with ordered signals are a special type of imperfect-information game. We can transform an ordered signal game into an imperfect-information game, and we will demonstrate this process. Given a game with ordered signals $\Gamma=(\tilde{\mathcal{N}},\tilde{H},\tilde{Z},\tilde{\rho},\tilde{A},\tilde{\chi},\tilde{\tau},\gamma,\Theta,\varsigma,O,\omega,\allowbreak\succeq,\tilde{u})$, we make some auxiliary definitions. Let $\mathfrak{r}=\{r=\gamma(\tilde{x})\mid\tilde{x}\in\tilde{X}\}$ be the set of reachable stages in this ordered signal game. We can partition the public nodes by stage, such that $\tilde{H}^{(r)}=\{h\in\tilde{H}\mid\gamma(h)=r\}$, $\tilde{Z}^{(r)}=\{z\in\tilde{Z}\mid\gamma(z)=r\}$, $\tilde{X}^{(r)}=\tilde{H}^{(r)}\cup\tilde{Z}^{(r)}$.

### v2

@@ 8 Experiment and Result
Experiments were conducted on Numeral211 Hold’em, a simplified version of two-player limit Texas Hold’em, which is divided into three phases; specific rules are detailed in Appendix D.2. In Numeral211 Hold’em, the quantity of abstracted signal infosets identified by each method at different phases is presented in Table 3. The performance of 0-Recall Outcome Isomorphism (i.e., Potential Outcome Isomorphism, POI), 2-Recall Outcome Isomorphism (2-ROI), and Lossless Isomorphism (LI) was evaluated in following experiments. Our experimental hardware are Linux servers with 96 cores, 2.40GHz of clock frequency, and 500GB of memory.
Phase 00-ROI $1$-ROI $2$-ROI LI
@@ 8.1 Asymmetric Abstraction in CFR Training
Our fisrt investigation focuses on the one-player-abstraction-perspective. Given a signal abstraction $\tilde{\alpha}^{\varTheta}=(\tilde{\alpha}^{\varTheta}_{1},\tilde{\alpha}^{\varTheta}_{2})$, we construct scenarios where player 1 employs the signal abstraction, while player 2 does not, resulting in the signal abstracted game $\tilde{\Gamma}^{1}=\tilde{\Gamma}^{(\tilde{\alpha}^{\varTheta}_{1},\varTheta_{2})}$. Conversely, we adjust the setup such that player 1 does not use abstraction, whereas player 2 does, thereby creating another signal abstracted game $\tilde{\Gamma}^{2}=\tilde{\Gamma}^{(\varTheta_{1},\tilde{\alpha}^{\varTheta}_{2})}$. Subsequently, by applying the chance-sampled counterfactual regret minimization (CSMCCFR) [38, 39], we iteratively solve for $\epsilon$-Nash equilibrium and map the resulting strategies back to the original game $\tilde{\Gamma}$ to obtain $\sigma^{1,*}=(\sigma^{1,*}_{1},\sigma^{1,*}_{2})$ and $\sigma^{2,*}=(\sigma^{2,*}_{1},\sigma^{2,*}_{2})$. Ultimately, by amalgamating these strategies into $\sigma^{*}=(\sigma^{1,*}_{1},\sigma^{2,*}_{2})$, which we called asymmetric abstraction Nash strategy, we adopt the $\sigma^{*}$ strategy in real-game scenarios as a means to counter strategies available in a non-abstracted setting from an abstracted viewpoint. The exploitability of $\sigma$ is then analyzed in terms of milli blinds per game (mb/g). The experiment involved $5.5\times 10^{10}$ iterations and took approximately 73 hours each to complete using two servers. The experimental results are depicted in Figure 6, where (a) showcases the results plotted on a linear scale of exploitability, while (b) presents the same results on a logarithmic scale.
@@ 8.2 Symmetric Abstraction in CFR Self-Play Training
Subsequently, we shift our focus to the setting of symmetric abstraction, which more closely aligns with the practical scenarios of AI training. Given a specific signal abstraction $\tilde{\alpha}^{\varTheta}$, we construct the corresponding signal abstracted game $\tilde{\Gamma}^{\tilde{\alpha}^{\varTheta}}$. Utilizing the CSMCCFR algorithm, we iteratively solve for the $\epsilon$-Nash equilibrium and map the resulting strategy back to the original game $\tilde{\Gamma}$, obtaining $\sigma^{*}$, and then calculate its exploitability. The experiment involved $5.5\times 10^{10}$ iterations and took approximately 106 hours to complete with a server. The experimental results are presented in Figure 7.
@@ Appendix C Perfect-Recall Games
    In any path of $\Gamma$, nodes from the same infoset $I\in\mathcal{I}_{i}$ of player $i$ appear at most once. In this case, for any node $x$ satisfying $I\sqsubseteq x$, $I[x]$ denotes the node in the path from $x^{o}$ to $x$ that belongs to $I$.
@@ D.1 Leduc Hold’em
Leduc Hold’em is a poker variant designed for two players, conceived primarily as a platform for computer game-playing research and introduced by Waugh et al. [40]. It was engineered to possess a strategic depth akin to Texas Hold’em while maintaining a manageable scale suitable for the development of intelligent gameplay strategies.
@@
    First Betting Phase: Following the distribution of hole cards, a phase of betting occurs. Players can choose to check or bet, with the bet size set at two chips.
@@
    Flop: After the initial betting phase, a single community card, termed the flop, is revealed from the deck.
@@
    Second Betting Phase: Another phase of betting takes place after the flop, with the bet size increasing to four chips.
@@
    Betting Options: Throughout the game, players have options to fold, call, or raise. Each betting phase permits only one bet and one raise per player, with fixed bet sizes of two chips in the first phase and four chips in the second.
@@ D.2 Numerall211 Hold’em
    First Betting Phase: Following the distribution of hole cards, a phase of betting occurs. Players can choose to check or bet, with the bet size set at 10 chips.
@@
    Flop: After the initial betting phase, a single community card, termed the flop, is revealed from the deck.
@@
    Second Betting Phase: Another phase of betting takes place after the flop, with the bet size increasing to 20 chips.
@@
    Turn: After the Second betting phase, another community card, termed the turn, is revealed from the deck.
@@
    Third Betting Phase: Another phase of betting takes place after the turn, with the bet size still set at 20 chips.
@@
    Betting Options: Throughout the game, players have options to fold, call, or raise. In each betting phase, the total sum of bets and raises is limited to a maximum of 4, with fixed bet sizes of 10 chips in the first phase and 20 chips in the last two betting phases.
@@ Appendix E A Transition from Games with Ordered Signals to Imperfect-Information Games
The games with ordered signals are a special type of imperfect-information game. We can transform an ordered signal game into an imperfect-information game, and we will demonstrate this process. Given a game with ordered signals $\Gamma=(\tilde{\mathcal{N}},\tilde{H},\tilde{Z},\tilde{\rho},\tilde{A},\tilde{\chi},\tilde{\tau},\gamma,\Theta,\varsigma,O,\omega,\allowbreak\succeq,\tilde{u})$, we make some auxiliary definitions. Let $\mathfrak{r}=\{r=\gamma(\tilde{x})\mid\tilde{x}\in\tilde{X}\}$ be the set of reachable phases in this ordered signal game. We can partition the public nodes by phase, such that $\tilde{H}^{(r)}=\{h\in\tilde{H}\mid\gamma(h)=r\}$, $\tilde{Z}^{(r)}=\{z\in\tilde{Z}\mid\gamma(z)=r\}$, $\tilde{X}^{(r)}=\tilde{H}^{(r)}\cup\tilde{Z}^{(r)}$.